OpenStudy (anonymous):

in the fraction 1/1/x = 1(x/1) = x ..... what rule allows this? in the fraction 1/1/x = 1(x/1) = x ..... what rule allows this? @Mathematics

OpenStudy (anonymous):

1/1/x = x/1 = x, multiplicative inverse or something like that. fractions of fractions. take the inverse of one fraction and multiply it by the other

OpenStudy (anonymous):

so, multiply the recip of the denominator by the numerator? or the recip of the whole thing? Sorry, I get confused when there are multiple fractions in a fraction

OpenStudy (anonymous):

most cases its multiply the numerator by the recip of the denom

OpenStudy (anonymous):

so x/1 x 1/1 would be the actual math.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wow. thanks, i cannot believe it took me this long to grasp that.

OpenStudy (anonymous):

1/2/4 = 1 * 4/2 = 1 * 2 = 2

OpenStudy (anonymous):

its less confusing to always use parenthesis... 1/(1/x) otherwise it could be (1/1)/x they are not the same

OpenStudy (anonymous):

i assume u know about inverse then?

OpenStudy (anonymous):

yeah, comes up a lot later in calculus

OpenStudy (anonymous):

1/a is the same as a^-1

OpenStudy (anonymous):

nevermind, i figured it out, but thanks for all your help.

OpenStudy (anonymous):

oh... haha i see they just switch ... is that true for all inverse where fog=x and gof=x?